OpenStudy (amenah8):

Find the CONCAVE DOWN and CONCAVE UP on the function f(x)=e^x - x^2 by setting the 2nd Derivative equal to 0!!

OpenStudy (amenah8):

First D= e^x - 2x Second D= e^x - 2 Set equal to 0--> e^x - 2=0 Now what??

zepdrix (zepdrix):

Hmm let's add 2 to each side maybe? :)

zepdrix (zepdrix):

\[\large\rm e^x=2\]Now we need to get that x out of the exponent some how... hmmm

zepdrix (zepdrix):

Remember anything about logs?

OpenStudy (amenah8):

a bit... so would we bring the x in front of the e?

zepdrix (zepdrix):

Log is the inverse operation of the exponential. When we take the composition of inverses, they essentially "undo" one another. \[\large\rm \ln(e^x)=x\ln(e)=x\] In the same way you would undo multiplication with its inverse (division), or undo subtraction with its inverse (addition, like we did earlier in this problem), you can undo exponentiation with its inverse (log).

zepdrix (zepdrix):

So we'll apply natural log to each side,\[\large\rm \ln(e^x)=\ln(2)\]

zepdrix (zepdrix):

\[\large\rm x=\ln(2)\]And that gives you your inflection point, ya? :)

OpenStudy (amenah8):

my inflection point?

OpenStudy (amenah8):

what is an inflection point? i am trying to find where the concave up and the concave down are located... i need to know if it is a positive or negative 2nd derivative

zepdrix (zepdrix):

Inflection point is where the function changes concavity :)

zepdrix (zepdrix):

Second derivative is ZERO at x=ln(2). That is the location where it changes from negative to positive! :)

myininaya (myininaya):

(ln(2),f(ln(2))) is a possible inflection point of f it is an inflection point if the concavity switches at that point you can choose test numbers from before and after ln(2) to check to see if (ln(2),f(ln(2)) is an inflection point (plug into f'' because that is what tells you if the function f is concave up/down) Concave up means f''>0 Concave down means f''<0 So you are testing the following two intervals: (-inf,ln(2)) and (ln(2),inf) by choosing a number from each interval and pluggin into f''

OpenStudy (solomonzelman):

Basically, \(\color{#000000 }{ \displaystyle f''(\ln 2 \pm \delta) }\) have to have different signs. where \(\delta\) is a sufficiently close value (like 0.3 or 1 even).

OpenStudy (solomonzelman):

have to have different signs, for \(x=ln(2)\) to be indeed the reflection point.

myininaya (myininaya):

:p inflection point

OpenStudy (solomonzelman):

Oh, lol, yes \(\color{red}{\rm in}\)flection point.

OpenStudy (solomonzelman):

Amenah ?

OpenStudy (amenah8):

sorry! i was doing other homework :)

OpenStudy (solomonzelman):

ok, so what can you say about this problem so far? Do you have questions about it, you understand it or.....

OpenStudy (amenah8):

how did you know the second derivative was 0 at x=ln(2)?

OpenStudy (solomonzelman):

You have found that the second derivative was, Example Problem: \(\color{#000000 }{ \displaystyle f''(x)=e^x-2 }\) Right

OpenStudy (amenah8):

right

OpenStudy (solomonzelman):

should have "example problem" there

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f''(x)=e^x-2 }\) you want to know at what value(s) of x is your second derivative is 0? So you set \(\color{#000000 }{ \displaystyle 0=e^x-2 }\)

OpenStudy (solomonzelman):

Then, you solve for x.

OpenStudy (amenah8):

and we did that by doing: ln(e^x) = ln2

OpenStudy (amenah8):

how would i use that to tell where the concave up/down is? is this right: x<ln(2) x>ln(2)

OpenStudy (amenah8):

how would i see if they are positive or negative?

OpenStudy (solomonzelman):

yes, plug in some values like ln(2)+0.1 and ln(2)-0.1 into the second derivative.

OpenStudy (solomonzelman):

If there are going to have different signs, then your cancavity indeed changes, and thus ln(2) is indeed an inflection point. If the sign is going to be the same, then x=ln(2) is not an inflection point.

OpenStudy (amenah8):

so i plug points into: e^x - 2x ?

OpenStudy (solomonzelman):

yes, your second derivative is \(\color{#000000 }{ \displaystyle f''(x)=e^x-2 }\) so you plug in \(\color{#000000 }{ \displaystyle x=\ln(2)\pm \delta }\) into the \(\color{#000000 }{ \displaystyle f''(x) }\) you can choose your delta to be \(\color{#000000 }{ \displaystyle \delta =1 }\)

OpenStudy (solomonzelman):

is this thing about delta too confusing?

OpenStudy (amenah8):

yeah, we haven't learned that yet :)

OpenStudy (solomonzelman):

well, I am just using that notation and name, all I want to know if the second derivative changes signs about x=ln(2)

OpenStudy (solomonzelman):

Do you want an example problem

OpenStudy (solomonzelman):

?

OpenStudy (amenah8):

sorry! i'm here :) yes, i would like an example, please.

OpenStudy (solomonzelman):

Find the concave down and concave up on the function \(\color{#000000 }{ \displaystyle f(x)=xe^x }\) by setting \(\color{#000000 }{ \displaystyle f''(x)=0 }\). We will differentiate this function through the product rule: \(\color{#000000 }{ \displaystyle d/dx~e^x=e^x }\) \(\color{#000000 }{ \displaystyle d/dx~x=1 }\) (So, \(\color{#000000 }{ \displaystyle d/dx~xe^x=e^x+xe^x }\)) Therefore, we know that, \(\color{#000000 }{ \displaystyle f'(x)=e^x+xe^x }\) And then we will differentiate this function again to get the second derivative. (note: we already know \(\color{#000000 }{ \displaystyle d/dx~xe^x }\)) \(\color{#000000 }{ \displaystyle f''(x)=e^x+(e^x+xe^x) }\) \(\color{#000000 }{ \displaystyle f''(x)=2e^x+xe^x }\) We want to find possible inflection points, so we set, \(\color{#000000 }{ \displaystyle 0=2e^x+xe^x }\) \(\color{#000000 }{ \displaystyle 0=e^x(2+x) }\) We know that \(\color{#000000 }{ \displaystyle e^x>0~\forall x\in\mathbb{R} }\), so dividing by \(e^x\) on both sides would have given us \(e^x=0\) as an alternative solution, but that is not a possibility since \(e^x\) is always positive for real numbers) \(\color{#000000 }{ \displaystyle 0=2+x}\) \(\color{#000000 }{ \displaystyle -2=x}\)

OpenStudy (solomonzelman):

Now, you would like to test \(x=2\) to see if this actually an inflection point!

OpenStudy (solomonzelman):

So, what you need to verify that on one sides of \(x=-2\), the second derivative is negative and on the other sides of \(x=-2\) is positive. (i.e., you want to verify the change in signs of the second derivative, which indicate the change in concavity of the function) review: the function is concave up at the x-values that satisfy f''(x)>0, the function concave down at the x values that satisfy f''(x)<0. Recall that, \(\color{#000000 }{ \displaystyle f''(x)=2e^x+xe^x }\) \(\color{#000000 }{ \displaystyle f''(-2+\color{blue}{1})=f''(-1)=2e^{-1}+(-1)e^{-1}=e^{-1}=\frac{1}{e}>0 }\) \(\color{#000000 }{ \displaystyle f''(-2-\color{blue}{1})=f''(-3)=2e^{-3}+(-3)e^{-3}=-e^{1}<0 }\) NOTE: I choose my delta; \(\color{#000000 }{ \displaystyle \delta =1 }\) (but choosing VERY large \(\delta\) might result in failure, however \(\delta=1\) is close enough) So we can imply that for \(\color{#000000 }{ \displaystyle x = (-\infty~,~ -2) }\), \(\color{#000000 }{ \displaystyle f''(x)<0 }\) AND for \(\color{#000000 }{ \displaystyle x = (-2~,~+\infty) }\), \(\color{#000000 }{ \displaystyle f''(x)>0 }\)

OpenStudy (solomonzelman):

So since the \(f''(x)\) indeed changes signs about \(x=-2\), (as I have shown choosing \(\delta=1\)), THEREFORE \(x=-2\) is an inflcetion point.

OpenStudy (solomonzelman):

And for (-2, ∞) the function is concave up (since second derivative is positive for these values), and for (-∞, -2) the function is caoncave down (since the second derivative is negative at those values)

OpenStudy (amenah8):

why is it -2? shouldn't it be positive?

OpenStudy (amenah8):

so in this problem, our inflection point is ln(2)

OpenStudy (solomonzelman):

in the example problem x=-2 is the inflection point. I showed how.

OpenStudy (amenah8):

oh! i see

OpenStudy (solomonzelman):

in oru problem the inflection point is ln(2), but you need to verify this, by evaluating \(f''(\ln(2)\pm\delta)\), and proving that they have different signs.

OpenStudy (solomonzelman):

what \(\delta\) do you want to choose? (has to be sufficiently small)

OpenStudy (amenah8):

1

OpenStudy (solomonzelman):

Ok, good, so you need to evaluate \(f''(\ln(2)\pm1)\)

OpenStudy (solomonzelman):

find f''(ln(2)+1) and f''(ln(2)-1)

OpenStudy (amenah8):

can you wait a moment while i grab my calculator?

OpenStudy (solomonzelman):

sure

OpenStudy (amenah8):

1.7 & -.31

OpenStudy (solomonzelman):

ok, so you know that for x>ln(2), the second derivative is positive, so the function is concave up for x>ln(2). ok, so you know that for x<ln(2), the second derivative is negative, so the function is concave down for x<ln(2).

OpenStudy (amenah8):

i got it! thank you so much :)

OpenStudy (solomonzelman):

You welcome!